How do you solve:
If four fifths + (- three tenths)= x + 1.5, then x = ?
Thank you

Answer by edjones(8007)

4/5 + (-3/10)=x+1.5
.8-.3=x+1.5 Change to decimals.
.5=x+1.5
x=-1 subtract -1.5 from each side
.
Ed